OpenStudy (anonymous):

Give the value of the discriminantof the equation 4x^-7x+3=0 a) 1 b)-128 c)0 d)128

OpenStudy (anonymous):

Answer is 1

OpenStudy (anonymous):

anna, do you know how to find the discriminant?

OpenStudy (anonymous):

Disc= b^2 - 4ac in this equation a=4, b=-7 and c=3 so, Disc= (-7)^2 - 4(4)(3) Disc= 49-48 Disc= 1

OpenStudy (anonymous):

@hsmt have you got it now..??

OpenStudy (anonymous):

@annawilliams40285 have you got it now..??

OpenStudy (anonymous):

yes, i had the answer quite some time ago. i was going to help her find the answer, not give it to her.